Write the molecular electronic configuration of N2+ molecule.

Grade:12

1 Answers

abhay gupta
3 years ago
The bond order of N2+ is 2.5 . Bond order = 1/2[Nb-Na], where, Nb=no. of electrons in bonding molecular orbital and Na= no. of electrons in anti-bonding molecular orbital.